OpenStudy (anonymous):

solve the formula for the perimeter of a rectangle,with width w and length y,for the length. p=2w+2l A) l=p-2w/2 B) l=p/2w -2 C) l=2w/p -2 D) l=p+2w/2

OpenStudy (triciaal):

width = w length = y= l p= 2l + 2w p = 2(l + w) divide each side by 2 p/2 = l + w subtract w from each side p/2 - w = l same as (p-2w)/2